OpenStudy (anonymous):

y= the square root of x^2-4....... how do i dertermine the domain and range

OpenStudy (anonymous):

\[y=\sqrt{x^2-4}\]

OpenStudy (anonymous):

for domain...under radical must be greater than or equal to zero or we can say it cant be negative thats the restriction we must consider for the case of radicals so\[x^2-4\ge0\]

OpenStudy (anonymous):

ok so 2 would be a possiblity for x

OpenStudy (anonymous):

do u know how to find all possible values for x from\[x^2-4\ge0\]?

OpenStudy (anonymous):

yeah but theres so many options even negative options

ganeshie8 (ganeshie8):

@sarah70 exactly! you need to see for what values of x, x^2-4 goes negative

ganeshie8 (ganeshie8):

and remove them from domain

ganeshie8 (ganeshie8):

test this : x is between 0 and 2 for ex, x = 1 : x^2-4 1^2-4 1-4 -3 so this 0<x < 2 should be removed from DOMAIN

OpenStudy (anonymous):

how do you know if x is between 0 and 2?

ganeshie8 (ganeshie8):

here, how i did : x^2-4 since 4 = 2^2, if i put x=2 above, it gives me 0 right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

you can see that anything more than 2 for x, makes x^2-4 a positive

ganeshie8 (ganeshie8):

does that make sense

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

so, [2, infiniti) is part of our domain, okay

ganeshie8 (ganeshie8):

we searched 0 to +infinity and we found those as domain. we need to still search 0 to -infiniti, okay ?

OpenStudy (anonymous):

isnt that also infinite?

ganeshie8 (ganeshie8):

you mean to say : (-infiniti, -2] ?

OpenStudy (anonymous):

huh? no i meant if you plug in all the negative values for x wouldnt the possiblities also be infinite?

ganeshie8 (ganeshie8):

oh ok thats right !! almost all -ve values of x, makes x^2-4 a positive, except few

ganeshie8 (ganeshie8):

test : x = -1

ganeshie8 (ganeshie8):

by now you should be able to see that : domain of \(\sqrt{x^2-4} \ is\ x \le -2, x \ge 2 \)

ganeshie8 (ganeshie8):

anything between, -2 and 2 makes the thing inside radical negative ! those values are not fit for domain

OpenStudy (anonymous):

but -1^2=1 and 1^2= 1 and theyre both greater than 0 so zero is the only # that doesnt apply so the domain is all real numbers except zero

ganeshie8 (ganeshie8):

but we need to see the *whole thing* inside radical : x^2-4

ganeshie8 (ganeshie8):

put x = -1, above. what you get ?

OpenStudy (anonymous):

oh NOW i understand

ganeshie8 (ganeshie8):

glad to hear :D

OpenStudy (anonymous):

but isnt there a certain way to write it?

ganeshie8 (ganeshie8):

yeah you can express them in "interval notation"

ganeshie8 (ganeshie8):

domain : (\(-\infty \), -2] U [2, \(\infty\))

OpenStudy (anonymous):

would i go about solving for the range the same way?

ganeshie8 (ganeshie8):

nopes

ganeshie8 (ganeshie8):

domain is about finding what all *input values* a function takes

ganeshie8 (ganeshie8):

range is about what all *output values* a function *gives out*

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

yup

ganeshie8 (ganeshie8):

so, just plugin the boundary values of domain, and see what the function gives out

ganeshie8 (ganeshie8):

we got the fomain as , \(x \ge 2\) : \(x \le -2\) :

ganeshie8 (ganeshie8):

domain*

ganeshie8 (ganeshie8):

put x = 2, 3, 4.... you can see that \(\sqrt{x^2-4}\) gives out, 0 and higher values

ganeshie8 (ganeshie8):

put x = -2, -3, -4... you can see that \(\sqrt{x^2-4}\) gives out, 0 and hight values AGAIN

ganeshie8 (ganeshie8):

higher*

ganeshie8 (ganeshie8):

do the range is \([0, \infty)\)

ganeshie8 (ganeshie8):

*so

OpenStudy (anonymous):

ok thanks for your time that was very detailed thank you :)

ganeshie8 (ganeshie8):

yw !! i did these very long time ago... i enjoyed going thru these again. thanks you too :)